OpenStudy (anonymous):

the areas of two circes are directly roportional to their radii squared (a/a1=r^2 / r1^2) Find the area A1 if A= 78.5, r=5 and r1=8.

OpenStudy (mertsj):

Is that supposed to be an equal sign instead of a plus sign?

OpenStudy (mertsj):

Then just plug in the given values and solve for a1

OpenStudy (anonymous):

200.96 yupz... simply put in the values

OpenStudy (phi):

you start with \[ \frac{A}{A_1} = \frac{r^2}{r_1^2}\] they tell you r is 5 and \(r_1 \) is 8, and A is 78.5. Replace the letters with these numbers. \[ \frac{78.5}{A_1} = \frac{5^2}{8^2}\] can you find \(A_1\) ? one way is "flip" both fractions and then multiply both sides by 78.5 and simplify another way is "cross multiply", then divide both sides by 5^2 (which is 25) can you finish ?

OpenStudy (anonymous):

yes,, i got it tyty!!

OpenStudy (phi):

yw